3 / x^2 - 25 = 1 / x + 5 + 2 / x - 5
Solve the equation for the variable

Answer by mananth(16946)

3 / (x^2 - 25) = 1 / (x + 5) + 2 / (x - 5)
..
It is advisable that the numerator and the denominator be separated by brackets.
..
3 / (x^2 - 25) = 1(x-5)+2(x+5) / (x^2-25)
...
3 / (x^2 - 25) = 1(x-5)+2(x+5) /(x^2-25)
..
3/(x^2-25)=x-5+2x+10/(x^2-25)
..
3/(x^2-25)= 3x+5/(x^2-25)
3=3x+5
3-5=3x
-2=3x
x=-2/3
..